OpenStudy (anonymous):

√28 simplify by factoring

OpenStudy (anonymous):

ok any idea how to start?

OpenStudy (nathan917):

2/7 i think

OpenStudy (anonymous):

sqrt 28 = sqrt (4*7) = 2* sqrt (7) = 2*2.646 = 5.29

OpenStudy (nathan917):

First we have to factor the radicand into square factors (Hint: *The best squares* try 100, 25, 16, 9 and 4. If none of these work, then try other square factors.) 4 is a factor of 28 so we can rewrite √28 as √4x7 Since 4 is a square factor, we take the square root of 4 which is 2. We put the 2 in front of the radical symbol: 2√7 Now, the radicand is 7 which does does not have square factor. Therefore the simplified radical is 2√7 . this is read "2 times the square root of 7" is how i got it.

OpenStudy (nathan917):

@Bev you sure?.......

OpenStudy (anonymous):

Yep! 4 and 7 are factors of 28, so 28 can be written as 4 times 7. That will simplify to 2 sqrt 7. By the way, how are you typing the square root sign. I don't see a square root sign on the keyboard. I am just abeginner, though. Thanks

OpenStudy (anonymous):

...

OpenStudy (nathan917):

Xd

OpenStudy (anonymous):

@Bev stop putting ur answer in decimals, those r approximations..

OpenStudy (anonymous):

no radicals are needed?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

it is 2 square root 7

OpenStudy (nathan917):

As i said........

OpenStudy (anonymous):

\[\sqrt{28}\]\[\sqrt{4}\sqrt{7}\]\[2\sqrt{7}\] that's ur answer, anyone who says other wise r idiots..

OpenStudy (nathan917):

@iforgot @HawkCrimson @Bev @higherlearning Case close :D

OpenStudy (nathan917):

* my answer ended this question :d *

OpenStudy (anonymous):

^ @nathan917 thx

OpenStudy (anonymous):

lol yu guys are too much

OpenStudy (nathan917):

:D he gets the medal.......

OpenStudy (anonymous):

lol this isnt a game to see who gets a medel. it's about teaching people :)

OpenStudy (nathan917):

Some what unfair but ah ill forget it i got 13 medals in less than 5 min..

OpenStudy (nathan917):

ik

OpenStudy (nathan917):

i just though my example was better. **i dont mean this in a bad way* :D

OpenStudy (anonymous):

@HawkCrimson " lol this isnt a game to see who gets a medel. it's about teaching people :)"

OpenStudy (anonymous):

i agree very much

OpenStudy (nathan917):

Same.

OpenStudy (nathan917):

K guys case closed dont spam on this question :d